A blue print is 1cm : 6 feet. The width of a building is 42 feet. What is the width of the building on the blueprint?
Answer by stanbon(75887)

A blue print is 1cm : 6 feet. The width of a building is 42 feet. What is the width of the building on the blueprint?
Use a proportion:
blue/bldg = blue/bldg
----
x/42 = 1/6
x = 42(1/6)
x = 7 cm
----------------
Cheers
Stan H.
================